Daily Events on 2024/07/02 (Tue)

last updated at  2024/07/02 10:34:13 JST


<< prev day

@

next day >>


Station

Pass No.
20240702-

Contact time (UT)

Max. Elev.
 (deg)

Events

AOS

LOS

KRN1

6020

07:17:45

07:31:10

50.6

OP/OG upload
EIS: TI-CMD
SOT: TI-CMD
XRT: TI-CMD

USC34

0100

08:36:29

08:48:43

24.3

OP/OG upload, XRT:TI-CMD, SOT:TI-CMD, EIS:TI-CMD (backup)
AOCS: Upload orbital element
AOCS: Upload pointing table
EIS: Table Upload
DR playback (by OP)

USC34

0200

10:13:30

10:25:04

22.9

DR playback (by OP)
XRT: Table Upload
SOT: Table Upload

USC34

0300

20:37:23

20:46:51

7.7

DR playback (by OP)

USC34

0400

22:12:49

22:26:24

69.2

DR playback (by OP)

Flare list(LM) : http://www.lmsal.com/solarsoft/last_events/


DR information

OP main: 2024/07/02 11:44:00 - 2024/07/04 10:57:00 (UT)

Remaining amount
(prediction)

2798 Mbits @ OP_START ( 2024/07/02 11:44:00 )


DR usage: The DR would become empty on Jul 3 at 6 UT if all telescopes took no images until that time.


SOT

XRT

EIS

DR allocation [%]

54

23

23

DR allocation [Mbits]

4024

1714

1714

DR TLMLMT [Mbits]

2113 @OP_START
-> 6215

2630

2630


daily note

Co-alignment program run:
Please run co-alignment program during the period for the three instruments.
- If the solar condition is not permitted, please delay to the next timeline.
- The sequence should start exactly 15 minutes after the pointing change to stabilize the satellite pointing drift.
- 2 hours for N-limb at *** 930*** arcsec and 2 hours for E limb at *** -970*** arcsec.
- Not necessary to be during SAA-free time.
- For XRT, take G-band data every 5 minutes instead of 8 minutes. This is just for the N limb run. 8 minutes cadence is still OK for the E limb run.
- For XRT, please take data even during SAA and twilight periods.
- For SOT, please use CTM HYSTERISIS (not CT RUN) at the start of each SP program
- For EIS, please run Imada's new co-alignment program

XRT:
- HOP409
- HOP484
- HOP448 (synoptics)
- HOP349 (4-6UT)
- AR obs

EIS:
- HOP409 (7/2 14 UT - 7/3 03 UT)
- HOP484 (SUNRISE-III)
- AR obs
- HOP130 was postponed.

SOT:
- HOP409
- HOP484 (SUNRISE-III)
- AR obs
- SP dark/flat


CHIEF
TOHBAN

SUB
TOHBAN

CP

SOT CO

XRT CO

EIS CO

EVENING

Kotanagi

Sakao

Okamoto/Natsume

Sakao

Itami

MORNING


operation files

pointing file : cmdpln/20240702/re-point_202407026020.txt